# Combining Deep Reinforcement Learning and Search with Generative Models for Game-Theoretic Opponent Modeling

### v2

@@ 2 Preliminaries
An $N$-player normal-form game consists of a set of players $\mathscr{N}=\{1,2,\dotsc,N\}$, $N$ finite pure strategy sets $\mathit{\Pi}_{i}$ (one per player) with a joint strategy set $\mathit{\Pi}=\mathit{\Pi}_{1}\times\mathit{\Pi}_{2}\times\cdots\mathit{\Pi}_{N}$, and a utility tensor (one per player), $u_{i}:\mathit{\Pi}\rightarrow\mathbb{R}$, and we denote player $i$’s utility as $u_{i}(\pi)$. Two-player (2P) normal-form games are called matrix games. A two-player zero-sum (purely adversarial) game is such that, $N=2$ and for all joint strategies $\pi\in\mathit{\Pi}:\sum_{i\in\mathscr{N}}u_{i}(\pi)=0$, whereas a common-payoff (purely cooperative) game: $\forall\pi\in\mathit{\Pi},\forall i,j\in\mathscr{N}:u_{i}(\pi)=u_{j}(\pi)$. A general-sum game is one without any restrictions on the utilities. A mixed strategy for player $i$ is a probability distribution over $\mathit{\Pi}_{i}$ denoted $\sigma_{i}\in\Delta(\mathit{\Pi}_{i})$, and a strategy profile $\sigma=\sigma_{1}\times\cdots\times\sigma_{N}$, and for convenience we denote $u_{i}(\sigma)=\mathbb{E}_{\pi\sim\sigma}[u_{i}(\pi)]$. By convention, $-i$ refers to player $i$’s opponents. A best response is a strategy $b_{i}(\sigma_{-i})\in\textsc{BR}(\sigma_{-i})\subseteq\Delta(\mathit{\Pi}_{i})$, that maximizes the utility against a specific opponent strategy: for example, $\sigma_{1}=b_{1}(\sigma_{-1})$ is a best response to $\sigma_{-1}$ if $u_{1}(\sigma_{1},\sigma_{-1})=\max_{\sigma_{1}^{\prime}}u_{1}(\sigma_{1}^{\prime},\sigma_{-1})$. An approximate $\epsilon$-Nash equilibrium is a profile $\sigma$ such that for all $i\in\mathscr{N},u_{i}(b_{i}(\sigma_{-i}),\sigma_{-i})-u_{i}(\sigma)\leq\epsilon$, with $\epsilon=0$ corresponding to an exact Nash equilibrium.
@@ 2.1 Combining MCTS and RL for Best Response
  function GenBR($i,\sigma,num\_eps$
@@
  function Search($s$, $\sigma$, $\bm{v},\bm{p},\bm{g}$
@@ 3 GenBR: Learning Best Response Search with a Generative Model
  function PSRO($\mathcal{G}$, MSS
@@ Appendix C Nash Bargaining Solution of Normal-form games via Projected Gradient Ascent
In this section, we elaborate on the method proposed in Section 4.2. As an abuse of notation, let $u_{i}$ denote the payoff tensor for player $i$ flattened into a vector; similarly, let $\mathbf{x}$ be a vector as well. Let $d$ be the number of possible joint strategies, e.g., $M^{N}$ for a game with $N$ agents, each with $M$ pure strategies. Let $\Delta^{d-1}$ denote the $(d-1)$-simplex, i.e., $\sum_{k=1}^{d}\mathbf{x}_{k}=1$ and $\mathbf{x}_{k}\geq 0$ for all $k\in\{1,\ldots,d\}$. We first make an assumption.
@@
Each outer product $u_{i}u_{i}^{\top}$ is positive semi-definite with eigenvalues $||u_{i}||$ (with multiplicity $1$) and 00 (with multiplicity $M^{N}-1$).
@@
$||\nabla f(\mathbf{x})||_{\infty}$ $=||\sum_{i}\frac{u_{i}}{u_{i}^{\top}\mathbf{x}-d_{i}}||_{\infty}$ (9)
$\leq\sum_{i}||\frac{u_{i}}{u_{i}^{\top}\mathbf{x}-d_{i}}||_{\infty}$ (10)
$=\sum_{i}\frac{||u_{i}||_{\infty}}{\gamma_{i}}$ (11)
@@
Let $\{\mathbf{x}^{t}\}$ be the sequence generated by Algorithm 4 with starting point $\mathbf{x}^{0}=d^{-1}\mathbf{1}$ and step size sequence $\alpha^{t}=\frac{\sqrt{(d-1)/d}}{||g^{\prime}(\mathbf{x}^{s})||_{2}}(t+1)^{-1/2}$. Then, for all $t>0$ one has
$\max_{\mathbf{x}\in\Delta^{d-1}}g(\mathbf{x})-\max_{0\leq s\leq t}g(\mathbf{x}^{s})$ $\leq\frac{\sqrt{2B_{\psi}(\mathbf{x}^{*},\mathbf{x}^{0})}||g^{\prime}(\mathbf{x}^{s})||_{2}}{\sqrt{t+1}}$ (13)
$\leq\frac{\sqrt{\frac{d-1}{d}}||g^{\prime}(\mathbf{x}^{s})||_{2}}{\sqrt{t+1}}$ (14)
$\leq\frac{\sqrt{d}||g^{\prime}(\mathbf{x}^{s})||_{\infty}}{\sqrt{t+1}}.$ (15)
where $g(x)=-f(x)$ is the log-Nash product defined in Sec 4.2.
@@
Given Assumption C.1, $f(x)$ is convex (Lemma C.2) and its gradients are bounded in norm (Lemma C.3). We then apply Theorem 4.2 of Beck and Teboulle [2003] with $f(x)=-g(x)$, $\psi(\mathbf{x})=\frac{1}{2}||\mathbf{x}||_{2}^{2}$ and note that $B_{\psi}(\mathbf{x}^{*},\mathbf{x}^{0})\leq\frac{1}{2}(d-1)/d$ to achieve the desired result. ∎
@@
Let $\{\mathbf{x}^{t}\}$ be the sequence generated by EMDA Beck and Teboulle [2003] with starting point $\mathbf{x}^{0}=d^{-1}\mathbf{1}$ and step size sequence $\alpha^{t}=\frac{\sqrt{2\log(d)}}{||g^{\prime}(\mathbf{x}^{s})||_{\infty}}(t+1)^{-1/2}$. Then, for all $t>0$ one has
$\max_{\mathbf{x}\in\mathcal{X}}g(\mathbf{x})-\max_{0\leq s\leq t}g(\mathbf{x}^{s})\leq\frac{\sqrt{2\log d}||g^{\prime}(\mathbf{x}^{s})||_{\infty}}{\sqrt{t+1}}$ (16)
where $g(x)=-f(x)$ is the log-Nash product defined in Sec 4.2.
@@
Given Assumption C.1, $f(x)$ is convex (Lemma C.2) and its gradients are bounded in norm (Lemma C.3). We then apply Theorem 5.1 of Beck and Teboulle [2003] with $f(x)=-g(x)$. ∎
@@ Appendix D Game Domain Descriptions and Details
Game $N$ Type Description From
@@
Kuhn poker (3P) 3 $N$P Zero-sum Lockhart et al. [2019]
Kuhn poker (4P) 4 $N$P Zero-sum Lockhart et al. [2019]
Leduc poker (3P) 3 $N$P Zero-sum Lockhart et al. [2019]
@@
Table 3: Benchmark games. $N$ is the number of players.
@@ D.1 Benchmark Games
    Each player is independently dealt one of $num$ items with uniform chance.
@@
    Player 1 makes one of $num$ utterances utterances, which is observed by player 2.
@@
    Player 2 makes one of $num$ utterances utterances, which is observed by player 1.
@@
    Both players privately request one of the $num$ items $\times$ $num$ items possible trades.
The trade is successful if and only if both player 1 asks to trade its item for player 2’s item and player 2 asks to trade its item for player 1’s item. Both players receive a reward of 1 if the trade is successful and 0 otherwise. We use $num$ items = $num$ utterances = 10.
@@
Sheriff is a simplified version of the Sheriff of Nottingham board game. The game models the interaction of two players: the Smuggler—who is trying to smuggle illegal items in their cargo– and the Sheriff– who is trying to stop the Smuggler. At the beginning of the game, the Smuggler secretly loads his cargo with $n\in\{0,...,n_{max}\}$ illegal items. At the end of the game, the Sheriff decides whether to inspect the cargo. If the Sheriff chooses to inspect the cargo and finds illegal goods, the Smuggler must pay a fine worth $p\cdot n$ to the Sheriff. On the other hand, the Sheriff has to compensate the Smuggler with a utility if no illegal goods are found. Finally, if the Sheriff decides not to inspect the cargo, the Smuggler’s utility is $v\cdot n$ whereas the Sheriff’s utility is 0. The game is made interesting by two additional elements (which are also present in the board game): bribery and bargaining. After the Smuggler has loaded the cargo and before the Sheriff chooses whether or not to inspect, they engage in r rounds of bargaining. At each round $i=1,...,r$, the Smuggler tries to tempt the Sheriff into not inspecting the cargo by proposing a bribe $b_{i}\in\{0,...b_{max}\}$, and the Sheriff responds whether or not they would accept the proposed bribe. Only the proposal and response from round r will be executed and have an impact on the final payoffs—that is, all but the $r^{th}$ round of bargaining are non-consequential and their purpose is for the two players to settle on a suitable bribe amount. If the Sheriff accepts bribe $b_{r}$, then the Smuggler gets $p\cdot n-b_{r}$, while the Sheriff gets $b_{r}$.
@@ Appendix E Hyper-parameters and Algorithm Settings
Hyperparameter Name Values in DoND Values in CT (PSRO w/ DQN BR) AlphaZero training episodes $num\_eps$ $10^{4}$ $10^{4}$ Network input representation an infoset vector $\in\mathbb{R}^{309}$ implemented in Lanctot et al. [2019] an infoset vector $\in\mathbb{R}^{463}$ Network size for policy & value nets [256, 256] for the shared torso [1024, 512, 256] Network optimizer SGD SGD UCT constant $c_{uct}$ 20 for IR and IE, 40 for SW, 100 for NP Returned policy type greedy towards $\bm{v}$ argmax over learned q-net # simulations in search 300 Learning rate for policy & value net 2e-3 for IR and IE, 1e-3 for SW, 5e-4 for NP 1e-3 Delayed # episodes for replacing $\bm{v},\bm{p},\bm{g}$ 200 200 Replay buffer size $\lvert D\rvert$ $2^{16}$ Batch size 64 128 PSRO empirical game entries # simulation 200 2000 L2 coefficient in evaluator $c_{1}$ 1e-3 Network size for generator net [300, 100] MLP Learning rate for generator net 1e-3 L2 coefficient in generator net $c_{2}$ 1e-3 PSRO minimum pure-strategy mass lower bound 0.005 0.005
@@ E.1 Generative Models
In DoND given an information state, the thing we only need to learn is the opponent’s hidden utilities. And since the utilities are integers from 0 to 10, we design our generative model as a supervised classification task. Specifically, we use the 309-dimensional state $s_{i}$ as input and output three heads. Each head corresponds to one item of the game. each head consists of 11 logits corresponding to each of 11 different utility values. Then each gradient step we sample a batch of $(s_{i},h)$ from the data buffer. And then do one step of gradient descent to minimize the cross-entropy loss $-h^{T}\log\bm{g}(s)+c_{2}\lVert\theta^{\prime}\rVert^{2}_{2}$, where here we overload $h$ to represent a one-hot encoding of the actual opponent utility vectors. And each time we need to generate a new state at information state $s_{i}$, we just feed forward $s_{i}$ and sample the utilities according to the output logits. But since we have a constraints $\mathbf{v}_{1}\cdot\mathbf{p}=\mathbf{v}_{2}\cdot\mathbf{p}=10$, the sampled utilities may not always be feasible under this constraint. Then we will do an additional L2 projection on to all feasible utility vector space and then get the final results.
@@ G.1 Overview of Approximate Nash equilibrium solving using PSRO
  • $N$-player Zero-Sum Games: Figure 8.
@@
Figure 8: Empirical Convergence to Nash Equilibria using Exact vs. DQN Best Response in $N$-Player Zero-Sum Benchmark Games.
@@ G.2 Deal or No Deal
For DQN and Boltzmann DQN, we used replay buffer sizes of $10^{5}$, $\epsilon$ decayed from $0.9\rightarrow 0.1$ over $10^{6}$ steps, a batch size of $128$, and swept over learning rates of $\{0.01,0.02,0.005\}$. For Boltzmann DQN, we varied the temperature $\eta\in{0.25,0.5,1}$. For all self-play policy gradient methods we used a batch size of 128, and swept over critic learning rate in $\{0.01,0.001\}$, policy learning rate in {.$\{.001,0.0005,0.0001\}$, number of updates to the critic before updating the policy in $\{1,4,8\}$, and entropy cost in $\{0.01,0.001\}$. DQN trained with the settings above and a learning rate of $0.005$ was the agent we found to achieve highest individual returns (and social welfare, and Nash bargaining score), so we select it as the representative agent for the independent RL category.
[…]
@@
[…]
@@
[…]
@@
[…]

### v3

@@ 2 Preliminaries
An $N$-player normal-form game consists of a set of players $\mathscr{N}=\{1,2,\dotsc,N\}$, $N$ finite pure strategy sets $\mathit{\Pi}_{i}$ (one per player) with a joint strategy set $\mathit{\Pi}=\mathit{\Pi}_{1}\times\mathit{\Pi}_{2}\times\cdots\mathit{\Pi}_{N}$, and a utility tensor (one per player), $u_{i}:\mathit{\Pi}\rightarrow\mathbb{R}$, and we denote player $i$’s utility as $u_{i}(\pi)$. Two-player (2P) normal-form games are called matrix games. A two-player zero-sum (purely adversarial) game is such that, $N=2$ and for all joint strategies $\pi\in\mathit{\Pi}:\sum_{i\in\mathscr{N}}u_{i}(\pi)=0$, whereas a common-payoff (purely cooperative) game: $\forall\pi\in\mathit{\Pi},\forall i,j\in\mathscr{N}:u_{i}(\pi)=u_{j}(\pi)$. A general-sum game is one without any restrictions on the utilities. A mixed strategy for player $i$ is a probability distribution over $\mathit{\Pi}_{i}$ denoted $\sigma_{i}\in\Delta(\mathit{\Pi}_{i})$, and a strategy profile $\sigma=\sigma_{1}\times\cdots\times\sigma_{N}$, and for convenience we denote $u_{i}(\sigma)=\mathbb{E}_{\pi\sim\sigma}[u_{i}(\pi)]$. By convention, i-i refers to player $i$’s opponents. A best response is a strategy $b_{i}(\sigma_{-i})\in\textsc{BR}(\sigma_{-i})\subseteq\Delta(\mathit{\Pi}_{i})$, that maximizes the utility against a specific opponent strategy: for example, $\sigma_{1}=b_{1}(\sigma_{-1})$ is a best response to $\sigma_{-1}$ if $u_{1}(\sigma_{1},\sigma_{-1})=\max_{\sigma_{1}^{\prime}}u_{1}(\sigma_{1}^{\prime},\sigma_{-1})$. An approximate $\epsilon$-Nash equilibrium is a profile $\sigma$ such that for all $i\in\mathscr{N},u_{i}(b_{i}(\sigma_{-i}),\sigma_{-i})-u_{i}(\sigma)\leq\epsilon$, with $\epsilon=0$ corresponding to an exact Nash equilibrium.
@@ 2.1 Combining MCTS and RL for Best Response
function GenBR($i,\sigma,num\_eps$)
@@
function Search($s$, $\sigma$, $\bm{v},\bm{p},\bm{g}$)
@@ 3 GenBR: Learning Best Response Search with a Generative Model
function PSRO($\mathcal{G}$, MSS)
@@ Appendix C Nash Bargaining Solution of Normal-form games via Projected Gradient Ascent
In this section, we elaborate on the method proposed in Section 4.2. As an abuse of notation, let $u_{i}$ denote the payoff tensor for player $i$ flattened into a vector; similarly, let $\mathbf{x}$ be a vector as well. Let $d$ be the number of possible joint strategies, e.g., $M^{N}$ for a game with $N$ agents, each with $M$ pure strategies. Let $\Delta^{d-1}$ denote the (d1)(d-1)-simplex, i.e., $\sum_{k=1}^{d}\mathbf{x}_{k}=1$ and $\mathbf{x}_{k}\geq 0$ for all $k\in\{1,\ldots,d\}$. We first make an assumption.
@@
Each outer product $u_{i}u_{i}^{\top}$ is positive semi-definite with eigenvalues $u_{i}||$ (with multiplicity $1$) and 0 (with multiplicity $M^{N}-1$).
@@
[…]
@@
[…]
where g(x)=f(x)g(x)=-f(x) is the log-Nash product defined in Sec 4.2.
@@
[…]
@@
[…]
where g(x)=f(x)g(x)=-f(x) is the log-Nash product defined in Sec 4.2.
@@
Given Assumption C.1, $f(x)$ is convex (Lemma C.2) and its gradients are bounded in norm (Lemma C.3). We then apply Theorem 5.1 of Beck and Teboulle [2003] with f(x)=g(x)f(x)=-g(x). ∎
@@ Appendix D Game Domain Descriptions and Details
Game NN Type Description From
@@
Kuhn poker (3P) 3 NNP Zero-sum Lockhart et al. [2019]
Kuhn poker (4P) 4 NNP Zero-sum Lockhart et al. [2019]
Leduc poker (3P) 3 NNP Zero-sum Lockhart et al. [2019]
@@
Table 3: Benchmark games. NN is the number of players.
@@ D.1 Benchmark Games
    Each player is independently dealt one of numnum items with uniform chance.
@@
    Player 1 makes one of numnum utterances utterances, which is observed by player 2.
@@
    Player 2 makes one of numnum utterances utterances, which is observed by player 1.
@@
    Both players privately request one of the numnum items $\times$ numnum items possible trades.
The trade is successful if and only if both player 1 asks to trade its item for player 2’s item and player 2 asks to trade its item for player 1’s item. Both players receive a reward of 1 if the trade is successful and 0 otherwise. We use numnum items = numnum utterances = 10.
@@
Sheriff is a simplified version of the Sheriff of Nottingham board game. The game models the interaction of two players: the Smuggler—who is trying to smuggle illegal items in their cargo– and the Sheriff– who is trying to stop the Smuggler. At the beginning of the game, the Smuggler secretly loads his cargo with $n\in\{0,...,n_{max}\}$ illegal items. At the end of the game, the Sheriff decides whether to inspect the cargo. If the Sheriff chooses to inspect the cargo and finds illegal goods, the Smuggler must pay a fine worth $p\cdot n$ to the Sheriff. On the other hand, the Sheriff has to compensate the Smuggler with a utility if no illegal goods are found. Finally, if the Sheriff decides not to inspect the cargo, the Smuggler’s utility is $v\cdot n$ whereas the Sheriff’s utility is 0. The game is made interesting by two additional elements (which are also present in the board game): bribery and bargaining. After the Smuggler has loaded the cargo and before the Sheriff chooses whether or not to inspect, they engage in r rounds of bargaining. At each round i=1,,ri=1,...,r, the Smuggler tries to tempt the Sheriff into not inspecting the cargo by proposing a bribe $b_{i}\in\{0,...b_{max}\}$, and the Sheriff responds whether or not they would accept the proposed bribe. Only the proposal and response from round r will be executed and have an impact on the final payoffs—that is, all but the $r^{th}$ round of bargaining are non-consequential and their purpose is for the two players to settle on a suitable bribe amount. If the Sheriff accepts bribe $b_{r}$, then the Smuggler gets $p\cdot n-b_{r}$, while the Sheriff gets $b_{r}$.
@@ Appendix E Hyper-parameters and Algorithm Settings
Hyperparameter Name Values in DoND Values in CT (PSRO w/ DQN BR) AlphaZero training episodes num_epsnum\_eps $10^{4}$ $10^{4}$ Network input representation an infoset vector $\in\mathbb{R}^{309}$ implemented in Lanctot et al. [2019] an infoset vector $\in\mathbb{R}^{463}$ Network size for policy & value nets [256, 256] for the shared torso [1024, 512, 256] Network optimizer SGD SGD UCT constant $c_{uct}$ 20 for IR and IE, 40 for SW, 100 for NP Returned policy type greedy towards $\bm{v}$ argmax over learned q-net # simulations in search 300 Learning rate for policy & value net 2e-3 for IR and IE, 1e-3 for SW, 5e-4 for NP 1e-3 Delayed # episodes for replacing $\bm{v},\bm{p},\bm{g}$ 200 200 Replay buffer size $\lvert D\rvert$ $2^{16}$ Batch size 64 128 PSRO empirical game entries # simulation 200 2000 L2 coefficient in evaluator $c_{1}$ 1e-3 Network size for generator net [300, 100] MLP Learning rate for generator net 1e-3 L2 coefficient in generator net $c_{2}$ 1e-3 PSRO minimum pure-strategy mass lower bound 0.005 0.005
@@ E.1 Generative Models
In DoND given an information state, the thing we only need to learn is the opponent’s hidden utilities. And since the utilities are integers from 0 to 10, we design our generative model as a supervised classification task. Specifically, we use the 309-dimensional state $s_{i}$ as input and output three heads. Each head corresponds to one item of the game. each head consists of 11 logits corresponding to each of 11 different utility values. Then each gradient step we sample a batch of $(s_{i},h)$ from the data buffer. And then do one step of gradient descent to minimize the cross-entropy loss $h^{T}\log\bm{g}(s)+c_{2}\lVert\theta^{\prime}\rVert^{2}_{2}$, where here we overload $h$ to represent a one-hot encoding of the actual opponent utility vectors. And each time we need to generate a new state at information state $s_{i}$, we just feed forward $s_{i}$ and sample the utilities according to the output logits. But since we have a constraints $\mathbf{v}_{1}\cdot\mathbf{p}=\mathbf{v}_{2}\cdot\mathbf{p}=10$, the sampled utilities may not always be feasible under this constraint. Then we will do an additional L2 projection on to all feasible utility vector space and then get the final results.
@@ G.1 Overview of Approximate Nash equilibrium solving using PSRO
  • NN-player Zero-Sum Games: Figure 8.
@@
Figure 8: Empirical Convergence to Nash Equilibria using Exact vs. DQN Best Response in NN-Player Zero-Sum Benchmark Games.
@@ G.2 Deal or No Deal
For DQN and Boltzmann DQN, we used replay buffer sizes of $10^{5}$, $\epsilon$ decayed from $0.9\rightarrow 0.1$ over $10^{6}$ steps, a batch size of $128$, and swept over learning rates of $\{0.01,0.02,0.005\}$. For Boltzmann DQN, we varied the temperature $\eta\in{0.25,0.5,1}$. For all self-play policy gradient methods we used a batch size of 128, and swept over critic learning rate in $\{0.01,0.001\}$, policy learning rate in $\{.001,0.0005,0.0001\}$, number of updates to the critic before updating the policy in $\{1,4,8\}$, and entropy cost in $\{0.01,0.001\}$. DQN trained with the settings above and a learning rate of $0.005$ was the agent we found to achieve highest individual returns (and social welfare, and Nash bargaining score), so we select it as the representative agent for the independent RL category.
Due the experimental limitation, we can only select 5 of our agents to human experiments. For convenience from now on we label a PSRO agent as $(\textit{MSS},\textit{BACKPROP\_TYPE},\textit{FINAL\_TYPE})$ if it is trained under meta-strategy solver MSS, its back-propagation type is BACKPROP_TYPE and we use FINAL_TYPE as its decision architecture. We apply standard empirical game theoretic analysis Wellman et al. [2025] on our agent pool: we create a 113x113 symmetric empirical game by simulating head to head results between every pair of our agents. During each simulation we toss a coin to assign the roles (first-mover or second-mover in DoND) to our agents. Then we decide the selected agents based on this empirical game matrix.
@@
Agent $\backslash$ Opponent IndRL Com1 Com2 Coop Fair
@@
Agent $\backslash$ Opponent IndRL Com1 Com2 Coop Fair
@@
Agent $\backslash$ Opponent IndRL Com1 Com2 Coop Fair